Assume a company's current ratio and acid-test ratio are less than 1.0 before it purchases inventory...

  1. Assume a company's current ratio and acid-test ratio are less than 1.0 before it purchases inventory on credit. When it makes the purchase:
    1. Its acid-test ratio decreases.                                  B) Its acid-test ratio remains unchanged.

C) Its current ratio decreases.                                    D) Its current ratio remains unchanged.

Please explain your answer. Thank you.
Answer #1

Answer A) Its acid-test ratio decreases.                                 

Suppose Current Assets are 80000 including 5000 of inventory and Current Liabilities are 100000 before it purchases inventory then

Current Ratio     = Current assets / Current Liabilities

                            = 80000 / 100000

                             = .80

Quick Ratio     = Quick assets / Current Liabilities

                            = 75000 / 100000

                             = .75

Suppose company purchases more inventory for 5000 on credit then after purchase of inventory on credit

Current Ratio     = Current assets / Current Liabilities

                          = 85000 / 105000

                             = .81

Quick Ratio     = Quick assets / Current Liabilities

                            = 70000 / 105000

                           = .67

Therefore, after purchase of inventory on credit its quick ratio will decrease.